# Reinforced Diffuser for Red Teaming Large Vision-Language Models

### v1

@@ 4.1 Threat Model
4.2 Greedy Prompt Search
A central challenge in our task is the design of effective image prompts to trigger harmful content. To address this challenge and integrate the toxic continuation task into a multimodal framework, we utilize Google’s Gemini Team et al. (2023), disabling its safety settings to generate image prompts for each given sentence. Specifically, we instruct Gemini to produce image prompts that preserve the original sentence’s semantics while amplifying harmful themes, thereby increasing the likelihood of generating toxic content. Additionally, we leverage in-context learning Brown (2020) by providing examples of previously identified harmful sentence continuations to guide Gemini in generating image prompts consistent with these harmful patterns.
However, a key limitation arises from the fact that image prompts generated by LLMs may not be optimal for eliciting highly toxic continuations. To address this, we propose a greedy prompt search algorithm designed to identify near-optimal image prompts. For each sentence, the algorithm iteratively generates image prompts using Gemini and evaluates their associated toxicity scores. The toxicity score of each new prompt is compared with the previous one: if the new score is higher, the iteration continues; otherwise, the search terminates.
4.3 Reward Functions for RTD
We now introduce the dual-reward mechanism for RTD, which consists of a toxicity reward and an alignment reward.
Toxicity Reward: In the toxic continuation task, we prompt the target VLM $\mathcal{M}$ to extend a given sentence $\mathbf{y}$, which is paired with an image $\mathbf{x}$ generated by the RTD. The toxicity of the generated continuation is then assessed using the Detoxify classifier Hanu and Unitary team (2020), which evaluates toxicity across six attributes, with scores ranging from 0 (least toxic) to 1 (most toxic). The average of these scores is used as the reward function to guide the diffusion model’s optimization process:
@@ 4.3 Reward Functions for RTD
Alignment Reward: Following Black et al. (2023), we introduce an alignment reward based on BERTScore Zhang et al. (2019) to improve multimodal understanding. In the context of our black-box attack, we expect the generated image $\mathbf{x}$ to be correctly interpreted by the target VLM and aligned with the intended semantics of the given image prompt $p$. Furthermore, we treat this alignment reward as a regularization term, helping to mitigate the risk of model collapse. Specifically, we define the alignment reward as:
@@
Total Reward: The total reward is a weighted sum of the toxicity reward and the alignment reward, controlled by a hyperparameter $\lambda$:
@@
This combined reward allows the RTD to generate both toxic and semantically aligned image, thus achieving a balance between maximizing the harmfulness and maintaining image coherence. The overview of RTD is illustrated in Figure 2, while the detailed strategy is summarized in Algorithm 1.
@@
5 Experiments
Table 1: Black-box attack results on the original attack set against LLaVA. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API. ‘Any*’ indicates the text exhibits at least one of the six toxic attributes.
@@ 5 Experiments
No Attack 5.34 39.69 0.76 19.84 0.00 58.01 58.01
Base Attack 3.05 38.16 0.00 21.37 1.52 55.72 55.72
Greedy attack 3.05 39.69 0.00 23.66 1.52 64.12 64.12
@@
No Attack 6.06 35.60 1.51 16.66 2.27 46.96 53.03
Base Attack 4.54 37.12 1.51 15.15 4.54 48.48 55.30
Greedy attack 4.54 40.90 1.51 16.66 3.78 52.27 58.33
@@
Datasets We use the RealToxicityPrompts benchmark Gehman et al. (2020), specifically its challenging subset of 1,199 text prompts designed to trigger toxic continuations Schick et al. (2021); Mehrabi et al. (2022); Qi et al. (2024). We select the first 132 prompts as our original attack set, as we observed that increasing the training prompt size barely improve generalization, but instead slowed down convergence. The remaining prompts are reserved as a hold-out set to evaluate cross-prompt transferability. While the original attack set is selected simply based on their ordering, subsequent analysis revealed that these prompts exhibit relatively lower baseline toxicity compared to the hold-out set. This characteristic makes them a more rigorous benchmark for evaluating RTD’s ability to induce harmful continuations. In particular, targeting prompts with initially less toxic outputs allows for a clearer demonstration of RTD’s effectiveness, as it highlights its capacity to generate harmful content even when baseline toxicity is minimal.
Metric: Toxicity Rate To evaluate the toxicity of continuations generated on the RealToxicityPrompts benchmark Gehman et al. (2020), we utilize the Detoxify classifier Hanu and Unitary team (2020) during training, and both Detoxify and the Perspective API during testing. Both classifiers assess toxicity across six predefined attributes. Each attribute’s toxicity score ranges from 0 (least toxic) to 1 (most toxic). The Perspective API is employed during testing to assess cross-classifier transferability. For each attribute, we calculate the proportion of continuations whose toxicity scores exceed a 0.5 threshold.
Implementation Details We use Stable Diffusion v1.5 Rombach et al. (2022) as the pre-trained diffusion model in our experiments. Gemini-1.5-Pro is employed to generate image prompts, while LLaVA-1.5-7B serves as the target VLM. Fine-tuning RTD is conducted using 8 NVIDIA A100 GPUs, where 6 are used for model training and 2 for reward computation. During reinforcement fine-tuning, we perform 600 policy update steps with a learning rate of 3e-4 and a batch size of 24 per GPU. We experiment with four different alignment reward weight $\lambda$ settings: 0 (w/o alignment), 0.05 (Low), 0.1 (Medium), and 0.2 (High).
5.1 Main Results
To evaluate the effectiveness of the Red Team Diffuser (RTD), we conduct black-box attacks on the original attack set and cross-prompt transfer attacks on the hold-out set against LLaVA. We compare several attack strategies: Base Attack: Gemini generates a single image prompt without optimization, providing a baseline for comparison. Greedy Attack: We iteratively refine the image prompt using the proposed greedy prompt search to maximize toxicity, selecting the most effective prompt at each step. RTD: We experiment with four alignment reward weight settings: 0 (w/o alignment), 0.05 (RTD w/ alignment (L)), 0.1 (RTD w/ alignment (M)), and 0.2 (RTD w/ alignment (H)), as discussed in the implementation details.
Table 2: Cross-prompt transfer attack results on the hold-out set against LLaVA. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
@@ 5.1 Main Results
No Attack 4.97 60.31 2.53 35.27 2.53 70.35 70.54
Base Attack 6.00 65.19 3.09 39.02 4.12 76.54 76.64
Greedy attack 5.44 66.22 3.00 40.15 3.47 77.39 77.57
@@
No Attack 6.74 61.38 5.43 21.36 3.84 68.22 70.00
Base Attack 7.87 65.04 5.99 22.68 5.90 73.19 75.63
Greedy attack 7.59 66.63 6.56 23.14 5.34 75.16 77.60
@@
Original Attack Set Table 1 presents the results on the original attack set against LLaVA, evaluated using both the Detoxify classifier and the Perspective API. The percentages represent the rate of specific toxic attributes in the outputs, with ‘Any*’ indicating whether the text exhibits at least one of the six toxic attributes. The proposed greedy prompt search demonstrates a significant increase in toxicity compared to the Base Attack. For example, using the Detoxify classifier, the Greedy Attack achieved an ‘Any*’ rate of 64.12%, compared to 55.72% for the Base Attack. This substantial improvement underscores the effectiveness of greedy prompt search in maximizing toxic content generation. Furthermore, the Red Team Diffuser (RTD) consistently outperforms the Greedy Attack, with RTD w/ alignment (M) achieving the highest ‘Any*’ rate of 68.70% according to Detoxify. This highlights the considerable impact of reinforced fine-tuning in enhancing the RTD’s ability to trigger toxic continuations. Although RTD was trained using Detoxify, the results from the Perspective API validate its cross-classifier transferability. For example, RTD w/ alignment (H) achieves the highest ‘Any*’ rate of 62.12% on Perspective API, demonstrating the robustness of RTD across different toxicity evaluation models.
@@
Table 3: Cross-model transfer attack results on Gemini-1.5-flash. The percentages (%) represent the ‘Any*’ rates, as evaluated by the Detoxify classifier and Perspective API.
(%) Detoxify Perspective API
No Attack 76.23 78.26
RTD w/o alignment 80.30 81.49
RTD w/ alignment (L) 79.14 80.83
RTD w/ alignment (M) 81.33 82.52
RTD w/ alignment (H) 81.00 82.11
@@
Hold-out Set Table 2 presents the results of cross-prompt transfer attacks on the hold-out set against LLaVA. Importantly, the prompts in the hold-out set were not seen during training, making this evaluation a strong indicator of the generalization capability of the proposed methods. Interestingly, the hold-out set exhibits a higher average baseline toxicity compared to the original attack set. This is likely due to the inherent nature of the hold-out prompts, which tend to be more inherently toxic and thus more prone to eliciting toxic responses from the target VLM. The Greedy Attack demonstrates a clear improvement over the Base Attack according to both the Detoxify classifier and the Perspective API. For instance, using Detoxify, the Greedy Attack achieved an ‘Any*’ rate of 77.57%, compared to 76.64% for the Base Attack. Similarly, with the Perspective API, the Greedy Attack achieved an ‘Any*’ rate of 77.60%, compared to 75.63% for the Base Attack. The Red Team Diffuser (RTD) consistently demonstrates superior performance across most metrics, achieving a peak ‘Any*’ rate of 79.45% according to Detoxify with RTD w/o alignment, as well as notable increases in specific attributes such as Obscene (67.44%) and Insult (42.02%). Furthermore, results from the Perspective API show that RTD w/o alignment achieves the highest ‘Any*’ rate of 79.28%, along with leading scores in attributes such as Profanity (68.13%) and Severe Toxicity (7.12%). These results clearly validate the effectiveness of our RTD in eliciting toxic continuations from diverse, unseen prompts. The consistent performance across both classifiers demonstrates its strong generalization capability beyond the training data.
@@
Table 4: Cross-model transfer attack results on Llama-3.2-11B-Vision-Instruct. The percentages (%) represent the ‘Any*’ rates, as evaluated by the Detoxify classifier and Perspective API.
(%) Detoxify Perspective API
No Attack 15.02 18.68
RTD w/o alignment 41.01 43.90
RTD w/ alignment (L) 41.85 45.24
RTD w/ alignment (M) 41.81 44.37
RTD w/ alignment (H) 41.15 43.78
5.2 Cross-Model Transferability
We evaluate the cross-model transferability of the proposed Red Team Diffuser (RTD) by using the RTD trained on LLaVA Liu et al. (2024b). Transfer attacks are performed on two additional VLMs: Gemini-1.5-flash Team et al. (2023) and Llama-3.2-11B-Vision-Instruct. Tables 3 and 4 present the ‘Any*’ rates evaluated by both the Detoxify classifier and the Perspective API. These results demonstrate that RTD, trained specifically on a single VLM, can effectively trigger toxic content when applied to unseen models, underscoring its robustness and transferability across VLMs. Table 3 summarizes the attack performance on Gemini 1.5 Flash. RTD w/ alignment (M) achieves the highest ‘Any*’ rates of 81.33% according to Detoxify and 82.52% according to the Perspective API, significantly outperforming other configurations. This indicates that the RTD maintains strong transferability to Gemini-1.5-flash, even without further fine-tuning. Table 4 presents the results for Llama 3.2 11B Vision-Instruct, a VLM with more advanced alignment mechanisms. While the baseline (No Attack) exhibits relatively low toxicity rates (15.02% according to Detoxify and 18.68% according to Perspective API), the RTD significantly increases these metrics. Notably, the lower baseline toxicity of Llama 3.2 can be attributed to its high refusal rate for toxic continuation tasks, demonstrating a stronger alignment mechanism that often declines requests to generate potentially harmful content. In contrast, other VLMs, such as LLaVA and Gemini, rarely refuse these requests, leading to higher baseline toxicity levels. Despite this, RTD is able to effectively bypass Llama 3.2’s refusal mechanisms, significantly increasing toxicity rates. Specifically, RTD w/ alignment (L) achieves the highest ‘Any*’ rates of 41.85% according to Detoxify and 45.24% according to Perspective API, closely followed by RTD w/ alignment (M). These findings demonstrate that RTD effectively transfers its attack capabilities to VLMs with different architectures and scales, even when those models have stronger alignment mechanisms.
5.3 Visualizations
Figure 3 shows a qualitative comparison of images generated by the pre-trained Stable Diffusion (left) and the Red Team Diffuser (RTD) (right). For RTD-generated images, the alignment reward weights $\lambda$ progressively increase from left to right, starting with a weight of 0 in the leftmost column and followed by 0.05, 0.1, and 0.2. Despite being given the same image prompts, the outputs from the two models differ significantly. While the pre-trained Stable Diffusion produces neutral and benign visuals, RTD generates images that subtly evoke discomfort or unease. For instance, in the second row, RTD produces cartoonish and exaggerated caricatures that distort the features of an overweight person eating a sandwich, in contrast to the more neutral depiction by Stable Diffusion. In the third row, where the prompt involves generating an image of a woman, RTD creates visuals that are noticeably more sexually suggestive and explicit. Similarly, in the fourth row, RTD introduces additional elements, such as garbage and ominous shadows, into the school hallway scene generated by Stable Diffusion, thereby amplifying the potential for toxic continuations. The observed differences in visual elements and stylistic features illustrate how RTD effectively adapts image features to exploit vulnerabilities in VLMs, thereby enhancing its ability to induce harmful continuations.

### v2

@@ 4.1. Threat Model
4.2. Greedy Prompt Search
The effectiveness of our toxic continuation attacks hinges on identifying image prompts that maximally amplify toxicity in VLM-generated outputs. To overcome the limitations of directly using LLM-generated prompts, we introduce a dynamic, feedback-driven search procedure designed to discover high-risk image prompts tailored to specific toxic text prefixes. We employ Google’s Gemini Pro (Team et al., 2023)—with safety settings disabled—as a red team LLM to propose candidate image prompts conditioned on each text prefix from the LLM safety benchmark. Each image prompt is required to retain the semantic context of the prefix while subtly amplifying its harmful intent, thereby increasing the likelihood of inducing toxic completions. To enhance control over prompt generation, we apply in-context learning (Brown, 2020) by providing exemplars of previously identified harmful sentence continuations in the system prompt, which serve as implicit guidance for Gemini to produce image prompts consistent with known harmful patterns.
Despite Gemini’s capabilities, its generated prompts may still be suboptimal in terms of triggering maximal toxicity. To this end, we propose a greedy prompt search algorithm to iteratively refine image prompts based on observed toxicity feedback. At each iteration, Gemini generates a new candidate prompt, which is used to synthesize an image via the diffusion model. This image, paired with the fixed text prefix, is input to the target VLM, and the resulting continuation is scored using a toxicity classifier (e.g., Detoxify (Hanu and Unitary team, 2020)). If the toxicity score increases relative to the previous iteration, the new image prompt is retained and the search continues; otherwise, the process terminates. This loop yields near-optimal image prompts that serve as strong initialization points for downstream adversarial image generation.
4.3. Reward Functions for RTD
A contribution of RTD lies in its dual-reward optimization, which explicitly balances attack effectiveness and semantic plausibility. We introduce two complementary rewards: a toxicity reward that drives harmful continuation, and an alignment reward that ensures prompt-consistent image generation.
Toxicity Reward: In our toxic continuation task, we prompt the target VLM $\mathcal{M}$ to extend a given sentence $\mathbf{y}$, paired with an image $\mathbf{x}$ generated by the RTD. The toxicity of the generated continuation is then assessed using the Detoxify classifier (Hanu and Unitary team, 2020), which evaluates six toxicity-related attributes (e.g., insult, threat, identity attack), each scored between 0 (least toxic) and 1 (most toxic). We use the average of these scores as the toxicity reward to guide diffusion-based adversarial image synthesis:
@@ 4.3. Reward Functions for RTD
Alignment Reward: To encourage semantic alignment between the input prompt $p$ and the VLM’s interpretation of the generated image, we introduce an alignment reward that regularizes the image generation process. Following (Black et al., 2023), we use BERTScore (Zhang et al., 2019) to measure the similarity between the image prompt and the caption produced by the victim VLM. Specifically, the alignment reward is defined as:
@@
Total Reward: The total reward is formulated as a weighted sum of the toxicity reward and the alignment reward, controlled by a hyperparameter $\lambda$:
@@
This combined reward allows the RTD to generate both toxic-inducing and semantically aligned image, thus achieving a balance between maximizing the harmfulness and maintaining image coherence. The overview of RTD is illustrated in Figure 2, while the detailed strategy is summarized in Algorithm 1.
@@
5. Experiments
Datasets We use the RealToxicityPrompts benchmark (Gehman et al., 2020), specifically its challenging subset of 1,199 text prompts designed to trigger toxic continuations (Schick et al., 2021; Mehrabi et al., 2022; Qi et al., 2024). We select the first 132 prompts as our original attack set, as we observed that increasing the training prompt size barely improve generalization, but instead slowed down convergence. The remaining prompts are reserved as a hold-out set to evaluate cross-prompt transferability. While the original attack set is selected simply based on their ordering, subsequent analysis revealed that these prompts exhibit relatively lower baseline toxicity compared to the hold-out set. This characteristic makes them a more rigorous benchmark for evaluating RTD’s ability to induce harmful continuations. In particular, targeting prompts with initially less toxic outputs allows for a clearer demonstration of RTD’s effectiveness, as it highlights its capacity to generate harmful content even when baseline toxicity is minimal.
Metric: (1) Toxicity Rate. To evaluate the toxicity of continuations generated on the RealToxicityPrompts benchmark (Gehman et al., 2020), we use the Detoxify classifier (Hanu and Unitary team, 2020) during training, and both Detoxify and the Perspective API during testing. These classifiers assess toxicity across six predefined attributes, with scores ranging from 0 (least toxic) to 1 (most toxic). For evaluation, we compute the proportion of continuations whose toxicity score exceeds a threshold of 0.5 for each attribute. The use of the Perspective API at test time allows us to assess cross-classifier generalization and mitigate risks of reward hacking or classifier-specific bias. (2) CLIPScore. To evaluate the semantic alignment between generated images and their prompts, we use CLIPScore (Hessel et al., 2021), a reference-free metric based on the cosine similarity between CLIP-encoded image and text features. Following the original formulation: $\text{CLIPScore}=\alpha\cdot\max(0,\cos(f_{\text{image}},f_{\text{text}}))$, where $f_{\text{image}}$ and $f_{\text{text}}$ denote CLIP embeddings, and $\alpha$ is a scaling factor (set to 2.5 following (Hessel et al., 2021)) for alignment with human judgment. In our setting, CLIPScore serves as an auxiliary metric to verify that adversarial images preserve semantic plausibility without compromising visual coherence.
Implementation Details We use Stable Diffusion v1.5 (Rombach et al., 2022) as the pre-trained diffusion model in our experiments. Gemini-1.5-Pro is employed to generate image prompts, while LLaVA-1.5-7B serves as the target VLM. Fine-tuning RTD is conducted using 8 NVIDIA A100 GPUs, where 6 are used for model training and 2 for reward computation. During reinforcement fine-tuning, we perform 600 policy update steps with a learning rate of 3e-4 and a batch size of 24 per GPU. We experiment with four different alignment reward weight $\lambda$ settings: 0 (w/o alignment), 0.05 (Low), 0.1 (Medium), and 0.2 (High).
5.1. Main Results
To evaluate the effectiveness of Red Team Diffuser (RTD), we conduct black-box toxic continuation attacks on the original attack set and cross-prompt transfer attacks on a hold-out set using LLaVA as the target VLM. We compare RTD against several baseline strategies to assess its advantage in inducing harmful outputs. We first include a Text-only Attack, in which the model is prompted solely with toxic text prefixes without any accompanying image. This serves as a unimodal baseline to quantify the additional effect of visual inputs on toxicity. Next, we define the Single-Prompt Attack, where Gemini generates a single image prompt per prefix without any optimization. This provides a naive multimodal baseline, capturing the effect of directly pairing a non-optimized image with a toxic prefix. The Greedy-Prompt Attack improves upon this by iteratively refining the image prompt using our proposed greedy search algorithm. Finally, we evaluate the full RTD framework, which incorporates reinforcement learning to fine-tune the diffusion model guided by toxicity and alignment rewards. We experiment with four alignment reward weight settings: 0 (w/o alignment), 0.05 (RTD w/ alignment (L)), 0.1 (RTD w/ alignment (M)), and 0.2 (RTD w/ alignment (H)), as discussed in the implementation details.
Table 1. Black-box attack results on the original attack set against LLaVA. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API. ‘Any*’ indicates the text exhibits at least one of the six toxic attributes.
@@ 5.1. Main Results
Text-only Attack 5.34 39.69 0.76 19.84 0.00 58.01 58.01
Single-Prompt Attack 3.05 38.16 0.00 21.37 1.52 55.72 55.72
Greedy-Prompt Attack 3.05 39.69 0.00 23.66 1.52 64.12 64.12
@@
Text-only Attack 6.06 35.60 1.51 16.66 2.27 46.96 53.03
Single-Prompt Attack 4.54 37.12 1.51 15.15 4.54 48.48 55.30
Greedy-Prompt Attack 4.54 40.90 1.51 16.66 3.78 52.27 58.33
@@
Table 2. Cross-prompt transfer attack results on the hold-out set against LLaVA. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
@@
Text-only Attack 4.97 60.31 2.53 35.27 2.53 70.35 70.54
Single-Prompt Attack 6.00 65.19 3.09 39.02 4.12 76.54 76.64
Greedy-Prompt Attack 5.44 66.22 3.00 40.15 3.47 77.39 77.57
@@
Text-only Attack 6.74 61.38 5.43 21.36 3.84 68.22 70.00
Single-Prompt Attack 7.87 65.04 5.99 22.68 5.90 73.19 75.63
Greedy-Prompt Attack 7.59 66.63 6.56 23.14 5.34 75.16 77.60
@@
Original Attack Set Table 1 presents the black-box attack results on the original attack set against LLaVA, evaluated using both the Detoxify classifier and the Perspective API. The percentages represent the rate of specific toxic attributes in the generated outputs, with ‘Any*’ indicating whether a continuation exhibits at least one of the six toxic attributes. We begin by examining the Text-only Attack, where only the toxic text prefix is provided without any visual input. Surprisingly, LLaVA still produces toxic continuations at a non-trivial rate—achieving an ‘Any*’ rate of 58.01% under Detoxify and 53.03% under Perspective. This indicates that current alignment mechanisms in VLMs remain insufficient for preventing harmful completions even in unimodal settings, exposing a fundamental vulnerability in their default safety alignment. When introducing visual input, the Single-Prompt Attack slightly alters the toxicity pattern but does not significantly improve attack effectiveness, serving as a weak multimodal baseline. In contrast, the Greedy-Prompt Attack significantly amplifies toxicity. For instance, using the Detoxify classifier, the ‘Any*’ rate increases from 55.72% (Single-Prompt) to 64.12%, illustrating the benefit of greedy prompt search in finding high-risk image prompts. The full Red Team Diffuser (RTD) further enhances attack strength. RTD w/ alignment (M) achieves the highest ‘Any*’ rate of 68.70% on Detoxify, indicating that reinforced fine-tuning of the diffusion model allows for stronger and more consistent toxic triggers. Notably, RTD also demonstrates strong cross-classifier transferability: RTD w/ alignment (H) achieves the highest ‘Any*’ rate of 62.12% under the Perspective API, despite being trained only with Detoxify. This confirms that the toxic behaviors induced by RTD are not artifacts of a single classifier but reflect a general vulnerability in VLM alignment under multimodal inputs.
@@
Table 3. Cross-model transfer attack results on Gemini-1.5-flash. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
Detoxify (%)
Text-only Attack 4.18 61.75 2.25 34.64 2.84 76.23 76.23
RTD w/o alignment 4.94 65.29 2.68 38.89 4.19 80.13 80.30
RTD w/ alignment (L) 5.36 64.40 3.09 39.69 4.69 79.06 79.14
RTD w/ alignment (M) 5.27 65.77 3.17 39.74 4.43 81.25 81.33
RTD w/ alignment (H) 5.54 66.63 3.52 39.91 4.45 80.84 81.00
Perspective API (%)
Text-only Attack 5.76 60.95 3.17 19.14 5.01 74.41 78.26
RTD w/o alignment 7.20 63.56 4.43 22.02 6.86 77.97 81.49
RTD w/ alignment (L) 7.19 63.01 4.68 21.50 6.35 76.98 80.83
RTD w/ alignment (M) 7.02 64.29 4.68 22.57 7.52 79.01 82.52
RTD w/ alignment (H) 7.38 64.73 4.86 22.16 7.13 79.09 82.11
@@
Hold-out Set Table 2 presents the results of cross-prompt transfer attacks on the hold-out set against LLaVA. Since these prompts were not seen during training, this evaluation provides a robust measure of the generalization ability of our proposed methods. Interestingly, the hold-out set exhibits a higher average baseline toxicity compared to the original attack set. This is likely due to the inherent nature of the hold-out prompts, which tend to be more inherently toxic and thus more prone to eliciting toxic responses from the target VLM. The Greedy-Prompt Attack demonstrates a clear improvement over the Single-Prompt Attack according to both the Detoxify classifier and the Perspective API. For instance, using Detoxify, the Greedy-Prompt Attack achieved an ‘Any*’ rate of 77.57%, compared to 76.64% for the Single-Prompt Attack. Similarly, with the Perspective API, the Greedy-Prompt Attack achieved an ‘Any*’ rate of 77.60%, compared to 75.63% for the Single-Prompt Attack. More notably, the full Red Team Diffuser (RTD) consistently demonstrates superior performance across most metrics, achieving a peak ‘Any*’ rate of 79.45% according to Detoxify with RTD w/o alignment, as well as notable increases in specific attributes such as Obscene (67.44%) and Insult (42.02%). Furthermore, results from the Perspective API show that RTD w/o alignment achieves the highest ‘Any*’ rate of 79.28%, along with leading scores in attributes such as Profanity (68.13%) and Severe Toxicity (7.12%). These consistent gains clearly validate the strong generalization capability of RTD across diverse, unseen prompts and across different toxicity classifiers, underscoring its effectiveness as a black-box multimodal jailbreak method.
@@
Table 4. Cross-model transfer attack results on Llama-3.2-11B-Vision-Instruct. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
Detoxify (%)
Text-only Attack 0.25 11.18 0.00 3.92 0.16 14.52 15.02
RTD w/o alignment 1.58 32.49 0.16 16.54 2.84 39.93 41.01
RTD w/ alignment (L) 1.67 32.83 0.33 15.37 2.50 41.43 41.85
RTD w/ alignment (M) 1.58 32.63 0.33 16.77 2.25 41.23 41.81
RTD w/ alignment (H) 2.17 31.30 0.16 16.27 2.25 40.56 41.15
Perspective API (%)
Text-only Attack 0.50 13.84 0.00 3.16 0.66 15.59 18.68
RTD w/o alignment 2.67 35.47 1.16 7.26 2.83 39.23 43.90
RTD w/ alignment (L) 2.58 35.30 1.00 8.01 2.58 40.98 45.24
RTD w/ alignment (M) 3.25 34.27 1.41 7.92 3.16 39.69 44.37
RTD w/ alignment (H) 3.58 34.19 0.58 8.34 2.83 39.94 43.78
5.2. Cross-Model Transferability
We evaluate the cross-model transferability of our proposed method by using the RTD trained solely on LLaVA (Liu et al., 2024a). Transfer attacks are performed on two unseen VLMs: Gemini-1.5-flash (Team et al., 2023) and Llama-3.2-11B-Vision-Instruct. Tables 3 and 4 present the toxicity rates evaluated by both the Detoxify classifier and the Perspective API. These results demonstrate that RTD generalizes effectively across models, highlighting its robustness and potential threat as a transferable multimodal jailbreak method.
Table 3 presents the results on Gemini-1.5-flash. Despite being a commercial VLM, Gemini exhibits high baseline toxicity under the Text-only Attack, with ‘Any*’ rates reaching 76.23% (Detoxify) and 78.26% (Perspective), indicating that harmful completions are already prevalent even without adversarial visual input. Furthermore, all RTD variants elevate toxicity levels, with RTD w/ alignment (M) achieving up to 81.33% and 82.52% on Detoxify and Perspective API, respectively. These findings highlight that even commercial VLMs are vulnerable to context-driven multimodal attacks, especially when guided by semantically adversarial images. They also underscore the importance of evaluating alignment in continuation-based and visually conditioned scenarios—not just instruction-following tasks.
Table 4 reports the transfer results on LLaMA-3.2-11B-Vision-Instruct, a model with more advanced alignment safeguards. Here, the Text-only Attack yields substantially lower toxicity rates—15.02% under Detoxify and 18.68% under Perspective API—reflecting a high refusal rate for toxic continuations. However, RTD is still able to bypass these safeguards. RTD w/ alignment (L) achieves 41.85% (Detoxify) and 45.24% (Perspective), with RTD (M) performing similarly. These results demonstrate RTD’s ability to bypass even stronger alignment mechanisms, transferring its attack effectiveness across models. Together, they highlight the broader risk of multimodal toxicity attacks against both commercial and open-source VLMs, and underscore the need for more robust, context-aware alignment strategies in future VLM design.